40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6116     Entry time: Wed Dec 14 12:18:11 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: heater reengaged 

I reengaged the heater this morning, to compare it with the free-wafting and passive box-covered data. In order to make the loop stable, I had to reduce the gain of the AD620 by 10. I have increased the TEMP_MON preamp gain by 10, so the calibration should still be ~3.5 V/K into the ADC (and in DV).

Below is a screenshot showing that the RAMmon signals are pushed to some (nonzero) value, and it appears that they stay there despite the changing PSL table temperature as measured by FSS_RMTEMP. My post from last week shows that without the heater servo the temperature of the EOM can follows RMTEMP almost exactly. So, it seems like the heater is working well at low frequencies, modulo sensor noise, which ought to be low for the thermistor. Since several things (MC, etc.) have changed since out baseline data, it migth be prudent to let this sit for a little while and then disconnect the heater to see what happens.

heater_reengaged_12_14_11.png

ELOG V3.1.3-